OpenStudy (anonymous):

a (b cos C – c cos B) = b^2 – c^2

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

what is your question?

OpenStudy (anonymous):

proving with law of sines and cosines

OpenStudy (anonymous):

well in triangle ABC we have \[b^2=a^2+c^2-2ac \cos B\]\[c^2=a^2+b^2-2ab \cos C\]now evaluate \(a (b \cos C – c \cos B)\) and see if that is equal to the right hand side of given equation

OpenStudy (anonymous):

when bracket is opened and multiplied and divided by 2

OpenStudy (anonymous):

we get part of it

OpenStudy (anonymous):

u can substitute \(\cos B\) and \(\cos C\) from equations i wrote using law of cosins

OpenStudy (anonymous):

oh so let me check...

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

\[a\left (b \frac{a^2+b^2-c^2}{2ab}-c \frac{a^2+c^2-b^2}{2ac} \right)=b^2-c^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i was stuck half way. when you gave the formula i got it

OpenStudy (anonymous):

when bracket is opened and multiplied by 2 we get it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome